OpenStudy (anonymous):

At what point does the graph f(x) = x^2e^-x^2 have a horizontal tangent line ?????? Have no idea :(

OpenStudy (anonymous):

What defines a horizontal tangent? The derivative (or slope) is 0, i.e. horizontal.

OpenStudy (anonymous):

Can you calculate the derivative of the function?

OpenStudy (anonymous):

LOL I can try just one moment

OpenStudy (anonymous):

\[f(x)=x^2e^{-x^2}\] Have I interpreted this correctly?

OpenStudy (anonymous):

yes that is right :)

OpenStudy (anonymous):

Ok. Any progress on the derivative?

OpenStudy (anonymous):

\[2x \times e ^{-x^2}+x^2 \times e ^{-x^2} \times -2x\]

OpenStudy (anonymous):

Yep and now to solve f'(x) = 0. Start with factorising.

OpenStudy (anonymous):

I have never factorised anything like this :(

OpenStudy (anonymous):

Start off by cleaning up. \[2xe^{-x^2}-2x^3e^{-x^2}\] That should help.

OpenStudy (anonymous):

Would I take out the common factor of \[2xe ^{-x^2}\]

OpenStudy (anonymous):

Yep that would work.

OpenStudy (anonymous):

\[2xe ^{-x^2} (1-x^2)\]

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

So now, f'(x)=0.

OpenStudy (anonymous):

0,1 ?

OpenStudy (anonymous):

Clearly the exponential function never equals 0, so that leaves x = 0 and 1 - x^2 = 0

OpenStudy (anonymous):

0, 1 and you missed an answer

OpenStudy (anonymous):

pardon?

OpenStudy (anonymous):

1 - x^2 = 0 has two solutions for x

OpenStudy (anonymous):

+ or - 1

OpenStudy (anonymous):

That's is. And those three solutions combined gives your answer.

OpenStudy (anonymous):

Excellent thank you so very much :)

OpenStudy (anonymous):

Your welcome!

OpenStudy (anonymous):

so do I say that the function will be horizontal when x = 0, 1 or -1?

OpenStudy (anonymous):

You do indeed.

OpenStudy (anonymous):

Thank you again :)